OpenStudy (mony01):

Find the area enclosed by parabolas x=y^2-4y, x=2y-y^2

ganeshie8 (ganeshie8):

start by drawing a rough sketch and find the intersecting points

OpenStudy (mony01):

to find the intersection points do i equal them to each other?

ganeshie8 (ganeshie8):

yes solve them simultaneously, but before even solving this, its a good to have some idea of how the area u wanted looks like..

ganeshie8 (ganeshie8):

clearly the given equations represent parabolas, facing east and west also if u think a bit, u wil notice that both will go through the point (0, 0) so, (0, 0) is one intersecting point. u still need to find other intersecting point by equating them...

OpenStudy (mony01):

how can i know how the graph looks like?

ganeshie8 (ganeshie8):

two ways :- 1) if u knw a bit about parabolas, : change the given equations to vertex form, then it becomes easy to visualize/sketch 2) simply enter the equations @ : https://www.desmos.com/calculator

ganeshie8 (ganeshie8):

pick one..

OpenStudy (mony01):

thanks i got the graph

OpenStudy (mony01):

so now i equal them to eachother?

ganeshie8 (ganeshie8):

yes

OpenStudy (mony01):

would the intervals be from 2 to 4?

ganeshie8 (ganeshie8):

nope

ganeshie8 (ganeshie8):

thats the area between parabolas u wanto find right ?

OpenStudy (mony01):

yea

ganeshie8 (ganeshie8):

intersecting points : (0, 0) and (-3, 3) so the area wud be : \(\large \mathbb{\int_0^3 (2y-y^2) - (y^2-4y) dy}\)

ganeshie8 (ganeshie8):

see if that makes more or less sense..

OpenStudy (mony01):

is the answer 27?

OpenStudy (mony01):

in wolframalpha do you how they simplify it to 6y-2y^2?

ganeshie8 (ganeshie8):

we have below : \(\large \mathbb{\int_0^3 (2y-y^2) - (y^2-4y) dy}\) group like terms : \(\large \mathbb{\int_0^3 (2y-y^2 - y^2+4y) dy}\) \(\large \mathbb{\int_0^3 (2y-2y^2+4y) dy}\) \(\large \mathbb{\int_0^3 (6y-2y^2) dy}\)

OpenStudy (mony01):

ok thank you for helping me